OpenStudy (anonymous):

What is the area of a sector with radius 10" and measure of arc equal to 45°?

OpenStudy (anonymous):

if the measure of the arc is 45 deg then the sector is 45/360 or 1/8 the area of the circle with radius 10. so just take the area of the circle and multiply it by 1/8: Sector Area = 1/8 * (area of circle) = 1/8 * (pi*radius^2) = 1/8 * 100pi = 100pi/8

OpenStudy (anonymous):

thnks here is another one i need help with